OpenStudy (anonymous):

The quotient of a number and five is eight less than the number. What is the number?

OpenStudy (anonymous):

Let the number be x x/5 = x-8 x = 5(x-8) x = 5x-40 4x = 40 x = 10

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np :)